[Koji, Yuta]
We found that C1:SUS-MC{1,2,3}_TO_COIL_3_4_GAIN was somehow changed to -1, and feedback signal for SIDE was fedback to LLCOIL, which is apparently not correct.
We checked the snapshots on May 25 and confirmed that it was used to be 0, so we fixed it.
We suspect that it happened during the beam spot measurement, because the measurement changes the TO_COIL matrix gains.
Now, we don't see any MC beam spot oscillation.
Quote: |
[Koji, Suresh, Jenne, Yuta]
Background:
We noticed that the beam spots on MC mirrors are oscillating in ~ 1 Hz yesterday. It means MC mirrors are actually oscillating. This was observable even if the WFS servo is off.
|
|